OpenStudy (anonymous):

For each of the following problems, set up an equation and solve. Write a sentence explaining your answer. 1. Renee wants to take her daughter and some of her daughter’s friends to the movies. A movie ticket is $8. Renee budgets $50. How many of her daughter’s friends can go to the movies?

OpenStudy (anonymous):

@aloud

OpenStudy (anonymous):

$8.00 div by $50.00= ?

OpenStudy (anonymous):

$6.25

OpenStudy (anonymous):

its 0.16

OpenStudy (anonymous):

i see i divided 50,00 in 8 first but then how many friends would she be able to take

OpenStudy (anonymous):

yup sorry so round it to the merest tenth

OpenStudy (anonymous):

2

OpenStudy (anonymous):

it would be 6 6 of her daughter and some of her daughter’s friends can go to the movie

OpenStudy (anonymous):

i ment 6 sorry

OpenStudy (anonymous):

i mean that was my answer at first but my teacher said she cannot take 6 friend bc renee and wanna go with her daughter to i don't get it

OpenStudy (anonymous):

its 2 against 1 who do you thank is wright

OpenStudy (anonymous):

i'm sorry she cannot take 6 friends to the movies do you want me to tell you why because i just found out on my own by looking at it again

OpenStudy (anonymous):

yeah go ahead

OpenStudy (anonymous):

Renee will have to pay $8 for each ticket. Since you do not know how many tickets this will be, you will have to use a variable, or letter to represent it. Let's say you use "t" for tickets. So Renee's receipt would look like: QTY x PRICE = TOTAL t 8 50 In algebra, we write the equation: 8t = 50 To figure out the number of tickets, you will have to divide by 8 on both sides (opposite math). This results in the answer: t = 6.25 Can Renee really buy 6.25 tickets? Is this possible? So what is the true answer for the word problem? How many tickets can she get for $50? 6 tickets, or 7 tickets? The answer would be 6 since it is not possible to buy part of a ticket. So, how many tickets does this leave available for her daughter's friends? 4 ,Because it is 6 tickets so Renee wanna take her daughter so 6-2=4 so she can only take 4 friends

OpenStudy (anonymous):

so i was right the first time when i said 8t divided by 50 =6.25

OpenStudy (anonymous):

yes good job you got it

OpenStudy (anonymous):

will you help me with 2 more

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

2. A plumber charges a flat rate of $50 and $25 an hour for the work. The plumber received $150 for his services. How many hours did the plumber work?

OpenStudy (anonymous):

50+25h=150 subtract 50 on both sides you will get 100 then divide 25 into 100 and get 4 The plumber only worked 4 hours. THIS IS WHAT I GOT

OpenStudy (anonymous):

WAIT don't delete it i was going to see how you did it because she said i did something wrong

OpenStudy (anonymous):

ok this one would also be 6 because 150 div by 25 =6

OpenStudy (anonymous):

Let's solve your equation step-by-step. 50f+25=150 Step 1: Subtract 25 from both sides. 50f+25−25=150−25 50f=125 Step 2: Divide both sides by 50. 50f50=12550 f=52 Answer: f=52

OpenStudy (anonymous):

its 6

OpenStudy (anonymous):

never mind she said my answer is right i put the answer as -8x when it should have said x=-8 but thanks for tryin to help me

OpenStudy (anonymous):

kk